OpenStudy (anonymous):

WILL FAN AND MEDAL PLZZZZ HELP!! A car travels 1/6 of the distance between two cities in 3/5 of an hour. At this rate, what fraction of the distance between the two cities can the car travel in 1 hour? THE ANSWER HAS TO BE A FRACTION!

OpenStudy (anonymous):

@KEYS

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

ANYONE PLZ HELP ME?

OpenStudy (afrodiddle):

Hold on.

OpenStudy (jdoe0001):

\(\large { \begin{array}{ccllll} distance&hour \\\hline\\ \frac{1}{6}&\frac{3}{5}\\ x&1 \end{array}\implies \cfrac{\frac{1}{6}}{x}=\cfrac{\frac{3}{5}}{1}\impliedby \textit{solve for "x"} }\)

OpenStudy (jdoe0001):

hint: cross-multiply

OpenStudy (anonymous):

3/30? @jdoe0001

OpenStudy (afrodiddle):

I'm afraid not.

OpenStudy (jdoe0001):

so.. what did you get for the cross-multiplication?

OpenStudy (anonymous):

3/30 @jdoe0001

OpenStudy (jdoe0001):

so \(\bf \cfrac{\frac{1}{6}}{x}=\cfrac{\frac{3}{5}}{1} \\ \quad \\ \cfrac{1}{6}\cdot 1=\cfrac{3}{5}\cdot x\) there... what would that give you for "x"?

OpenStudy (anonymous):

Idk? @jdoe0001

OpenStudy (jdoe0001):

well... you may want to brush up your fraction multiplication first I'd think

OpenStudy (anonymous):

What will the answer be? @jdoe0001 Is 3/30 wrong?

OpenStudy (jdoe0001):

well.. is not that

OpenStudy (anonymous):

ohh then I don't get it it's a little confusing? @jdoe0001

OpenStudy (jdoe0001):

\(\bf \cfrac{\frac{1}{6}}{x}=\cfrac{\frac{3}{5}}{1} \implies \cfrac{1}{6}\cdot 1=\cfrac{3}{5}\cdot x \implies \cfrac{1}{6}=\cfrac{3}{5}x\implies \cfrac{1}{6}=\cfrac{3x}{5}\) if you cross-multiply that, and solve for "x".... what would that give?

OpenStudy (anonymous):

??????? @jdoe0001 I don't remember learning that

OpenStudy (jdoe0001):

hmmm wondering how are you meant to solve it then is a direct variation.. .thus why I set it up as a proportion

OpenStudy (anonymous):

\[\frac{5}{18} \]

OpenStudy (anonymous):

Is that right @robtobey

OpenStudy (anonymous):

\[\frac{1}{6}=\frac{3 x}{5} \] Multiply each side by 5/3. The result will be that x has a cooeficient of "1", one.

OpenStudy (anonymous):

ohhhhh thanks :) can u help me with one more problem? @robtobey

OpenStudy (anonymous):

Ok. By the way, coefficient was misspelled in my posting above.

OpenStudy (anonymous):

Ellen wants to purchase a book that is regularly priced at $18. The book is discounted 15%. She also needs to pay a 6% sales tax on the discounted price. What is the total amount Ellen will pay for the book?

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

Is the answer 16.22 that's what I got? @robtobey

OpenStudy (anonymous):

Take 106% of the result of 85% times 18 for the total amount.

OpenStudy (anonymous):

Is 16.22 wrong? @robtobey

OpenStudy (anonymous):

That is what I calculated when the final calculation was rounded to the nearest cent. Some stores will probably always round up to the next cent.

OpenStudy (anonymous):

So that mean its 16.218 your answer? @robtobey

OpenStudy (anonymous):

16.218, not yet rounded, is what the Mathematica program delivered.

OpenStudy (anonymous):

So what's the answer then? @robtobey

OpenStudy (anonymous):

My guess is 16.22, which is 16.218 rounded to the nearest cent.

OpenStudy (anonymous):

Okay I will put 16.22 but thanks for your time and help! :) @robtobey

OpenStudy (anonymous):

Your welcome and thank you for the medal.

OpenStudy (anonymous):

Np @robtobey